OpenStudy (anonymous):

Look at the expression below. 2h + y / 9h^2 - y^2 - 4h^2 / 3h + y a. (3h + y) b. (3h - y)(3h + y) c. (3h + y)(3h + y)(3h - y) d. 3(h + y)(h - y) help me please :)

jimthompson5910 (jim_thompson5910):

it's this right? \[\Large \frac{2h + y}{9h^2 - y^2} - \frac{4h^2}{3h + y}\]

jimthompson5910 (jim_thompson5910):

just want to be sure I interpreted the problem correctly

OpenStudy (anonymous):

okay sorry was using the restroom :P

jimthompson5910 (jim_thompson5910):

that's ok

OpenStudy (anonymous):

yeahs that right.

jimthompson5910 (jim_thompson5910):

You factor the first denominator to get \[\Large \frac{2h + y}{9h^2 - y^2} - \frac{4h^2}{3h + y}\] \[\Large \frac{2h + y}{(3h - y)(3h + y)} - \frac{4h^2}{3h + y}\] The factored denominators are (3h - y)(3h + y) and (3h + y) The unique factors in the denominator are (3h - y) and (3h + y) So that means the LCD is (3h - y)(3h + y). You just multiply those unique factors you find in the denominator.

jimthompson5910 (jim_thompson5910):

I factored 9h^2 - y^2 by using the difference of squares rule a^2 - b^2 = (a - b)(a + b)

OpenStudy (anonymous):

ffs i can't learn this pellet omfg ima die -_-

jimthompson5910 (jim_thompson5910):

just take it one step at a time

OpenStudy (anonymous):

i am but i can't seem to get it right :(

jimthompson5910 (jim_thompson5910):

it's ok, you'll get it eventually

OpenStudy (anonymous):

:;/ let me see.

OpenStudy (anonymous):

well D is the only that is different, so i do not think its that 1.

OpenStudy (anonymous):

am i wrong?

jimthompson5910 (jim_thompson5910):

I wrote the answer above lol

jimthompson5910 (jim_thompson5910):

you're looking for the LCD right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i didn't see that xD

jimthompson5910 (jim_thompson5910):

that's ok

OpenStudy (anonymous):

okay so its B? lol

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

you're welcome